``Before, solving the problem, I just want to give you the heads up. It's
Systems of Equation: Application. I can't find the specific title to relate
with my story problems. My chapter in the book is Chapter Eight, Solving Systems by Substiution, Additional. Hope it helps!
*The Perimeter of a rectangle is 42 inches. If the length is doubled and the width is tripled, the perimeter becomes 102 inches. Find the length and width of the original rectangle.*

Answer by Mathtut(3670)

42=2L+2W--->2L=42-2W.....eq 1
:
102=2(2L)+3(2W)......eq 2
:
take 2L's value from eq 1 and plug it into eq 2
:
102=2(42-2W)+6W
:
102=84-4W+6W
:
18=2W
:
W=9 in
:
L=(42-2(9))/2=24/2=12 in